OpenStudy (anonymous):

A badminton court has an area of 880 feet. the width of the court is 20 feet. what is the length of the badminton court?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

sure! \[Area = Length \times Height\] do you know how to go from here?

OpenStudy (anonymous):

btw if you didn't know badminton courts are rectangular.

OpenStudy (anonymous):

so is it 880 X 20 ?

OpenStudy (anonymous):

well the area is 880 right? And the width is 20 so substitute those values for their corresponding label.

OpenStudy (anonymous):

what do you mean by corresponding label?

OpenStudy (anonymous):

in the formula so let me do "area" for you. So the area of the court is 880 so if we substitute it in the formula, the formula becomes: \[800 = Length \times Height\]

OpenStudy (anonymous):

sorry I mean \[880 = Length \times Height\]

OpenStudy (anonymous):

ok so 880= 20 x ? thats where im lost

OpenStudy (anonymous):

yup so we got: \[880 = 20 \times Length\] now we need to "solve for the length" we do this by isolating whatever we are looking for by itself.

OpenStudy (anonymous):

Since we have 20 times Length, to isolate Length we simply do the opposite function to both sides. Since it's 20 and multiplication, we use division so we should divide 20 on both sides. Does this make sense?

OpenStudy (anonymous):

so 880 x 20 x20?

OpenStudy (anonymous):

How is that dividing on both sides? Take it step by step. We need to divide 20 on each side. \[880 \div 20 = 20 \times Length \div 20\] Do you see how I have divided 20 on both sides?

OpenStudy (anonymous):

btw always always write out the whole equation with the equal sign, if you are solving for something. This is very important!

OpenStudy (anonymous):

is the nswer 44?

OpenStudy (anonymous):

before I answer that does it make sense what I did?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good! yes that is the answer as the equation becomes: \[44 = Length\] because 880/20 is 44.

OpenStudy (anonymous):

ok thank you can i ask you one more?

OpenStudy (anonymous):

yes sure

OpenStudy (anonymous):

you are playing the game four square on a 12 foot by 12 foot court. your square is 6 foot by 6 foot. what is the area of the four square court not including your square?

OpenStudy (anonymous):

find the area of the big square and subtract the small square. Do you know to do so?

OpenStudy (anonymous):

so 12x6

OpenStudy (anonymous):

whats the area of the big square?

OpenStudy (anonymous):

144?

OpenStudy (anonymous):

good! so whats the area of the small square?

OpenStudy (anonymous):

36

OpenStudy (anonymous):

so whats the difference between the two

OpenStudy (anonymous):

112

OpenStudy (anonymous):

check your math

OpenStudy (anonymous):

so basically subtracting?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

108 is the answer

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

np